On planet Tehar, the free-fall acceleration is the same as that on Earth but there is also a strong downward electric field that is uniform close to the planet's surface. A 2.00 kg ball having a charge of 6.00 µC is thrown upward at a speed of 23.8 m/s and it hits the ground after an interval of 4.40 s. What is the potential difference between the starting point and the top point of the trajectory?
Answer #1

Using force balance on ball

Fnet = -mg - qE = ma

(1). E = m*(-a - g)/q

Now using the equation in vertical motion for full flight

dy = vi*t + 0.5*a*t^2

dy = 0 = vi*t + 0.5*a*t^2

(2). a = -2*vi/t

for the upward part of flight, vf = 0

vi^2 + 2*a*h = vf^2 = 0

vi^2 - 2*2*vi*h/t = 0

h = vi*t/4

Now using these values in eq 1 and 2

E = (m/q)*(2*vi/t - g)

Now we know that

V = electric potential = E*h

V = (m/q)*(2*vi/t - g)*(vi*t/4)

V = (2*10^6/6)*(2*23.8/4.4 - 9.81)*(23.8*4.4/4)

V = 8798066.67